OpenStudy (anonymous):

All of the following expressions simplify to -4, except _____. -1 - 3 5 - (-9) 2(-2) -8 + 4

OpenStudy (anonymous):

well: -1 - 3 = -1 + (-3) = -4 5 - (-9) = 5 + 9 = 14 2(-2) = -4 -8 + 4 = -4

OpenStudy (anonymous):

5-(-9)

OpenStudy (anonymous):

as it gives 14

OpenStudy (anonymous):

got it @Bookreader15 ?

OpenStudy (anonymous):

get it bookreader

OpenStudy (anonymous):

@Bookreader15 medal mee

OpenStudy (anonymous):

plzz

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

ok,doesn't matter :) as @openstudyneedsme helped too , i'll give him a medal.